# thanks: Corresponding Author

### v1

@@ V Decays to strange mesons
$\widehat{\Pi}_{\mu\nu}^{1}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{I_{\mu}^{B_{s}^{\ast}}(y)$ (58)
$\times I^{D_{s}}(0)I_{\nu}^{\dagger}(x)\}|0\rangle,$
@@
$\langle 0|I_{\mu}^{B^{\ast}}|B_{s}^{\ast 0}(p^{\prime},\varepsilon)\rangle=f_{B_{s}^{\ast}}m_{B_{s}^{\ast}}\varepsilon_{\mu},$
$\langle 0|I^{D_{s}}|D_{s}^{+}(q)\rangle=\frac{f_{D_{s}}m_{D_{s}}^{2}}{m_{c}+m_{s}}.$ (60)
@@
Figure 4: SR data for $G_{1}(Q^{2})$ and $G_{2}(Q^{2})$, and extrapolating functions $\widehat{\mathcal{Z}}_{1}(Q^{2})$ (solid curve), and $\widehat{\mathcal{Z}}_{2}(Q^{2})$ (dot-dashed curve). The points $Q^{2}=-m_{\overline{D}_{s}^{\ast}}^{2}$ and $Q^{2}=-m_{D_{s}}^{2}$ are denoted by the rectangle and triangle, respectively.
@@
$\widehat{\Pi}_{\mu\nu}^{1\mathrm{OPE}}(p,p^{\prime})$ $=$ $\frac{1}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle\overline{c}c\rangle\mathrm{Tr}\left[\gamma_{\mu}S_{s}^{ij}(y)\right$. (61)
$\left.\times\gamma_{5}S_{c}^{ja}(-x)\gamma_{\nu}\gamma_{5}S_{b}^{ai}(x-y)\right].$
@@
$\widehat{\Pi}_{\mu\nu}^{2}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{I_{\mu}^{B_{s}}(y)$ (65)
$\times I^{D_{s}^{\ast}}(0)I_{\nu}^{\dagger}(x)\}|0\rangle,$
@@
$\widehat{\Pi}_{\mu\nu}^{2\mathrm{Phys}}(p,p^{\prime})=\frac{G_{2}(q^{2})\Lambda f_{B_{s}}m_{B_{s}}^{2}f_{D_{s}^{\ast}}m_{D_{s}^{\ast}}}{(m_{b}+m_{s})\left(p^{2}-m^{2}\right)(p^{\prime 2}-m_{B_{s}}^{2})}$
$\times\frac{1}{(q^{2}-m_{D_{s}^{\ast}}^{2})}\left[\frac{m^{2}+m_{B_{s}}^{2}-q^{2}}{2}g_{\mu\nu}-\frac{m^{2}}{m_{D_{s}^{\ast}}^{2}}p_{\mu}^{\prime}p_{\nu}^{\prime}\right$.
$\left.+\frac{m^{2}-m_{D_{s}^{\ast}}^{2}}{m_{D_{s}^{\ast}}^{2}}p_{\mu}^{\prime}p_{\nu}-\frac{m^{2}+m_{B_{s}}^{2}-q^{2}}{2m_{D_{s}^{\ast}}^{2}}(p_{\mu}p_{\nu}+p_{\mu}p_{\nu}^{\prime})\right]$
$+\cdots.$ (67)
@@
$\langle 0|I^{B_{s}}|B_{s}^{\ast 0}(p^{\prime})\rangle=\frac{f_{B_{s}}m_{B_{s}}^{2}}{m_{b}+m_{s}}$
$\langle 0|I_{\mu}^{D_{s}^{\ast}}|D_{s}^{\ast+}(q,\varepsilon)\rangle=f_{D_{s}^{\ast}}m_{D_{s}^{\ast}}\varepsilon_{\mu},$ (68)
@@
$\langle B_{s}^{0}(p^{\prime})D_{s}^{\ast+}(q,\varepsilon)|\mathfrak{M}(p,\epsilon)\rangle=G_{2}(q^{2})\left[(p\cdot p^{\prime})(\epsilon\cdot\varepsilon^{\ast})\right$.
$\left.-(p^{\prime}\cdot\epsilon)(p\cdot\varepsilon^{\ast})\right].$ (69)
@@
$\widehat{\Pi}_{\mu\nu}^{2\mathrm{OPE}}(p,p^{\prime})$ $=$ $\frac{1}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle\overline{c}c\rangle\mathrm{Tr}\left[\gamma_{5}S_{s}^{ij}(y)\right$. (70)
$\left.\times\gamma_{\mu}S_{c}^{ja}(-x)\gamma_{\nu}\gamma_{5}S_{b}^{ai}(x-y)\right].$
@@ VI Discussion and concluding remarks
The mass $m=(9740\pm 70)~\mathrm{MeV}$ has been extracted using the SR method with a rather high accuracy and may be confronted with available predictions obtained by employing alternative approaches. The result for $m$ permits us to conclude that $J/\psi B_{c}^{+}$ and $\eta_{c}B_{c}^{\ast+}$ lie above corresponding two-meson thresholds and easily breaks down to constituent mesons. Even in the lower value, the mass $m=9660~\mathrm{MeV}$ is still considerably above $9372~\mathrm{MeV}$ which makes these decays kinematically possible. In other words, the molecules $J/\psi B_{c}^{+}$ and $\eta_{c}B_{c}^{\ast+}$ are resonant structures and do not form bound states. In this aspect they differ from the axial-vector molecules $\Upsilon B_{c}^{-}$ and $\eta_{b}B_{c}^{\ast-}$ which at the low-mass scenario are bound states.
@@
Hadronic molecules built of only heavy quarks were on focus of studies in Ref. Liu:2024pio . The authors used the extended local gauge formalism to explore such molecules. They did not find axial-vector bound states in $cc\overline{c}\overline{b}$ sector, which agree with our conclusions. But they did not give information on a mass(es) of possible resonant structures. In the lack of such calculations for the hadronic molecules, it is interesting to compare our results with predictions made for relevant diquark-antidiquark states.
The diquark-antidiquark exotic mesons with different spin-parities and nonsymmetrical quark contents were studied in Refs. An:2022qpt ; Galkin:2023wox (and references therein). The constituent quark model was used to evaluate parameters of the fully heavy tetraquarks in Ref. An:2022qpt , whereas the relativistic quark model employed in the second paper. The mass of the axial-vector tetraquark $cc\overline{c}\overline{b}$ in these articles were found equal to $9706~\mathrm{MeV}$and $9611~\mathrm{MeV}$, respectively. First, one sees a $100~\mathrm{MeV}$ discrepancy in predictions of different methods. Nevertheless, we fix the mass gaps around of $34~\mathrm{MeV}$ and $130~\mathrm{MeV}$ between diquark-antidiquark and molecular states of the same content, which are natural for particles with different internal organizations: Colored diquarks and antidiquarks form more tightly bound compounds than ordinary colorless mesons.

### v2

@@ V Decays to strange mesons
$\widehat{\Pi}_{\mu\nu}^{1}(p,p^{\prime})=i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{I_{\mu}^{B_{s}^{\ast}}(y)I^{D_{s}}(0)I_{\nu}^{\dagger}(x)\}|0\rangle,$ (60)
@@
$\langle 0|I_{\mu}^{B^{\ast}}|B_{s}^{\ast 0}(p^{\prime},\varepsilon)\rangle=f_{B_{s}^{\ast}}m_{B_{s}^{\ast}}\varepsilon_{\mu},\ \ \ \langle 0|I^{D_{s}}|D_{s}^{+}(q)\rangle=\frac{f_{D_{s}}m_{D_{s}}^{2}}{m_{c}+m_{s}}.$ (62)
@@
Figure 4: SR data for $G_{1}(Q^{2})$ and $G_{2}(Q^{2})$, and extrapolating functions $\widehat{\mathcal{Z}}_{1}(Q^{2})$ (solid curve), and $\widehat{\mathcal{Z}}_{2}(Q^{2})$ (dot-dashed curve). Values of these functions at the points $Q^{2}=-m_{\overline{D}_{s}^{\ast}}^{2}$ and $Q^{2}=-m_{D_{s}}^{2}$ are shown by the rectangle and triangle, respectively.
@@
$\widehat{\Pi}_{\mu\nu}^{1\mathrm{OPE}}(p,p^{\prime})=\frac{1}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle\overline{c}c\rangle\mathrm{Tr}\left[\gamma_{\mu}S_{s}^{ij}(y)\gamma_{5}S_{c}^{ja}(-x)\gamma_{\nu}\gamma_{5}S_{b}^{ai}(x-y)\right].$ (63)
@@
$\widehat{\Pi}_{\mu\nu}^{2}(p,p^{\prime})=i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{I_{\mu}^{B_{s}}(y)I^{D_{s}^{\ast}}(0)I_{\nu}^{\dagger}(x)\}|0\rangle,$ (67)
@@
$\widehat{\Pi}_{\mu\nu}^{2\mathrm{Phys}}(p,p^{\prime})$ $=$ $\frac{G_{2}(q^{2})\Lambda f_{B_{s}}m_{B_{s}}^{2}f_{D_{s}^{\ast}}m_{D_{s}^{\ast}}}{(m_{b}+m_{s})\left(p^{2}-m^{2}\right)(p^{\prime 2}-m_{B_{s}}^{2})}\frac{1}{(q^{2}-m_{D_{s}^{\ast}}^{2})}\left[\frac{m^{2}+m_{B_{s}}^{2}-q^{2}}{2}g_{\mu\nu}-\frac{m^{2}}{m_{D_{s}^{\ast}}^{2}}p_{\mu}^{\prime}p_{\nu}^{\prime}\right$. (69)
$\left.+\frac{m^{2}-m_{D_{s}^{\ast}}^{2}}{m_{D_{s}^{\ast}}^{2}}p_{\mu}^{\prime}p_{\nu}-\frac{m^{2}+m_{B_{s}}^{2}-q^{2}}{2m_{D_{s}^{\ast}}^{2}}(p_{\mu}p_{\nu}+p_{\mu}p_{\nu}^{\prime})\right]+\cdots.$
@@
$\langle 0|I^{B_{s}}|B_{s}^{\ast 0}(p^{\prime})\rangle=\frac{f_{B_{s}}m_{B_{s}}^{2}}{m_{b}+m_{s}},\ \ \langle 0|I_{\mu}^{D_{s}^{\ast}}|D_{s}^{\ast+}(q,\varepsilon)\rangle=f_{D_{s}^{\ast}}m_{D_{s}^{\ast}}\varepsilon_{\mu},$ (70)
@@
$\langle B_{s}^{0}(p^{\prime})D_{s}^{\ast+}(q,\varepsilon)|\mathfrak{M}(p,\epsilon)\rangle=G_{2}(q^{2})\left[(p\cdot p^{\prime})(\epsilon\cdot\varepsilon^{\ast})-(p^{\prime}\cdot\epsilon)(p\cdot\varepsilon^{\ast})\right].$ (71)
@@
$\widehat{\Pi}_{\mu\nu}^{2\mathrm{OPE}}(p,p^{\prime})=\frac{1}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle\overline{c}c\rangle\mathrm{Tr}\left[\gamma_{5}S_{s}^{ij}(y)\gamma_{\mu}S_{c}^{ja}(-x)\gamma_{\nu}\gamma_{5}S_{b}^{ai}(x-y)\right].$ (72)
@@ VI Discussion and concluding remarks
The mass $m=(9740\pm 70)~\mathrm{MeV}$ has been extracted using the SR method with a rather high accuracy which means that the mass of the hadronic molecule is inside of the region $m\in[9660,9810]~\mathrm{MeV}$. This result permits us to conclude that $J/\psi B_{c}^{+}$ and $\eta_{c}B_{c}^{\ast+}$ lie above corresponding two-meson thresholds and easily breaks down to constituent mesons, because even in its lower value the mass $m=9660~\mathrm{MeV}$ is considerably above the $9372~\mathrm{MeV}$ threshold. In other words, the molecules $J/\psi B_{c}^{+}$ and $\eta_{c}B_{c}^{\ast+}$ are resonant structures and do not form bound states. Nevertheless, we treat $J/\psi B_{c}^{+}$ and $\eta_{c}B_{c}^{\ast+}$ as highly unstable hadronic molecules because they are formed by two heavy mesons and their interpolating currents are constructed in accordance with these structures.
It is instructive to compare this situation with the axial-vector molecule(s) $\mathcal{M}_{\mathrm{AV}}=\Upsilon B_{c}^{-}$ [and $\eta_{b}B_{c}^{\ast-}$] built of $bb\overline{b}\overline{c}$ quarks. It was demonstrated that the mass of $\mathcal{M}_{\mathrm{AV}}$ resides in the region $15710,15890]~\mathrm{MeV}$, and in the lower border of this interval $15710~\mathrm{MeV}$ it is stable against decays to mesons $\Upsilon B_{c}^{-}$ and $\eta_{b}B_{c}^{\ast-}$. The reason is that this mass is smaller than corresponding kinematical thresholds $15735~\mathrm{MeV}$ and $15737~\mathrm{MeV}$ Agaev:2025wyf . In this sense the mesons $\Upsilon$ and $B_{c}^{-}$, as well as $\eta_{b}$ and $B_{c}^{\ast-}$ may form bound-state axial-vector molecules. But even in this case they transform to ordinary mesons due to $\overline{b}b$ annihilations in these structures.
In other words, the axial-vector state $\Upsilon B_{c}^{-}$ is more stable against strong decays (in lower mass limit) than the molecule $J/\psi B_{c}^{+}$. A conclusion about stable nature of some of four-quark mesons containing $bb$ quarks is not new for the tetraquark spectroscopy. Thus, it was found that axial-vector diquark-antidiquark state $T_{bb}^{-}=bb\overline{u}\overline{d}$ lies below the $B^{-}\overline{B}^{\ast 0}$ and $B^{-}\overline{B}^{0}\gamma$ thresholds and is stable against the strong and radiative transformations Karliner:2017qjm ; Eichten:2017ffp ; Agaev:2018khe . Hence, it decays to conventional mesons via weak processes considered in Ref. Agaev:2018khe . The charmed counterpart of $T_{bb}^{-}$, i.e., a diquark-antidiquark state $T_{cc}^{+}=cc\overline{u}\overline{d}$ is above a $D^{\ast}D$ threshold and is not a stable particle Navarra:2007yw . It is remarkable that $T_{cc}^{+}$ was observed by the LHCb collaboration in $D^{0}D^{0}\pi^{+}$ mass distribution as a narrow peak with the width $\Gamma=(48\pm 2^{+0}_{-14})~\mathrm{keV}$ Aaij:2021vvq ; LHCb:2021auc . The mass of $T_{cc}^{+}$ is very close to the $D^{0}D^{\ast+}$ threshold $3875.1~\mathrm{MeV}$, but is smaller than this limit. This discovery confirmed experimentally that the axial-vector state $T_{cc}^{+}$ is strong-interaction unstable structure. The resonance $T_{cc}^{+}$ were investigated in the context of diquark and hadronic molecule models in numerous publications (for details see, Refs. Agaev:2021vur ; Agaev:2022ast ; Ling:2021bir ; Feijoo:2021ppq and references therein).
@@
Information on the partial widths of different decay channels as well as full width of the molecule $\mathfrak{M}$ are important for experimental analysis of these particles. A broad peak in the mass distributions of final-state $J/\psi B_{c}^{+}$ or $\eta_{c}B_{c}^{\ast+}$ mesons in some processes would indicate on observation of all-heavy axial-vector exotic meson with quark content $cc\overline{c}\overline{b}$. But a diquark-antidiquark state $cc\overline{c}\overline{b}$ also decays through these channels. Therefore, a conclusion about internal structure of such a peak can only be made by comparing experimental results with theoretical predictions for widths of the molecule and diquark-antidiquark states. Similarly, some enhancements may be seen in the mass distributions, for instance, $B^{\ast+}D^{0}$ and $B^{\ast 0}D^{+}$ mesons. But to interpret them as product of subdominant processes, one first needs to compute parameters of hadronic molecules $B^{\ast+}D^{0}$ and $B^{\ast 0}D^{+}$ which may dissociate to these mesons as well. Because exotic hadronic molecules with asymmetric contents are new objects of studies these problems, in particular decays of such states, are waiting for their analyses.
Hadronic molecules built of only heavy quarks were on focus of studies in Ref. Liu:2024pio . The authors used the extended local gauge formalism to explore such molecules. They did not find axial-vector bound states in $cc\overline{c}\overline{b}$ sector, which agree with our conclusions. But they did not give information on a mass(es) of possible resonant structures. In the lack of such calculations, it is interesting to compare our results with predictions made for relevant diquark-antidiquark states.
The diquark-antidiquark exotic mesons with different spin-parities and nonsymmetrical quark contents were studied in Refs. Liu:2019zuc ; Gordillo:2020sgc ; Deng:2020iqw ; Mutuk:2022nkw ; Zhang:2022qtp ; An:2022qpt ; Galkin:2023wox . In these articles the authors used different methods and approaches such as nonrelativistic constituent quark, multiquark color flux-tube, and chromomagnetic models, the Bethe-Salpeter equations and effective potential model to investigate fully heavy diquark-antidiquark structures. Consider, for instance last two publications. The constituent quark model was used to evaluate parameters of the fully heavy tetraquarks in Ref. An:2022qpt , whereas the relativistic quark model employed in the second paper Galkin:2023wox . The mass of the axial-vector tetraquark $cc\overline{c}\overline{b}$ in these articles were found equal to $9706~\mathrm{MeV}$ and $9611~\mathrm{MeV}$, respectively. First, one sees a $100~\mathrm{MeV}$ discrepancy in predictions of different methods. Nevertheless, we fix the mass gaps around of $34~\mathrm{MeV}$ and $130~\mathrm{MeV}$ between diquark-antidiquark and molecular states of the same content, which are natural for particles with different internal organizations: Colored diquarks and antidiquarks form more tightly bound compounds than ordinary colorless mesons.
The molecules $J/\psi B_{c}^{+}$ and $\eta_{c}B_{c}^{\ast+}$ are only building units of the physical states with asymmetric quark content. Real particles may be formed owing to mixing of the molecules $\mathfrak{M}$ and $\widetilde{\mathfrak{M}}$. Transitions between molecules $J/\psi B_{c}^{+}$ and $\eta_{c}B_{c}^{\ast+}$ as well as their mixing with axial-vector tetraquark $cc\overline{c}\overline{b}$ are among possible scenarios. But related problems require additional detailed studies which are beyond of the current article.